# Privacy vs. Profit: The Impact of Google’s Manifest Version 3 (MV3) Update on Ad Blocker Effectiveness

### v1

@@ 5.3 Selection of Websites for the Browser-Based Experiment
For the main sample of our browser-based experiment, we selected ad-supported websites to observe the impact of ad blockers and evaluate their effectiveness. For this purpose, we utilized a curated list from HestiaLabs [^5] https://github.com/hestiaAI/ad-radar, which comprised 195 websites. We selected these websites for their reliance on ad-supported content and their use of the Prebid framework, guaranteeing compatibility with the Azerion Ad Expert extension (see Section 5.2). This compatibility is crucial for ensuring the extension accurately records the number of ads, especially for the baseline browser instance using no ad blocker.
After manually reviewing each website to ensure the ad blockers and Azerion Ad Expert operated as expected, we refined the list to 114 websites. We excluded websites that redirected users, used ad-blocking prevention mechanisms, or were inaccessible. The final list for our main sample represents 114 popular, ad-supported websites across 10 countries.
We also conducted the browser-based experiment on two additional stratified samples for robustness tests. We used BuiltWith [^6] https://builtwith.com data of websites relying on the Prebid framework to conduct the stratified sampling: we stratified one sample by website employee count, yielding 191 websites, and another by Tranco [^7] https://tranco-list.eu popularity rank [48], yielding 185 websites. These stratified samples captured varying levels of organizational resources and popularity and are more representative of the overall population of ad-supported websites, enhancing the generalizability of our findings (see Section 6.4 and Appendix B.1).
@@ 5.4 Definition of Two Metrics for Ad Blocker Effectiveness
For the first metric, the number of blocked ads, we used the extension Azerion Ad Expert. As mentioned (see Section 5.2), this tool recorded the number of display ad slots on a website. By subtracting the baseline number of ads on the website from the measured number with the ad blocker, we obtained the measure of the number of blocked ads.
To validate the accuracy of these ad counts, we manually counted ads in collected screenshots from a random subsample of 100 website–ad-blocker observations. A Pearson correlation analysis showed a strong positive correlation ($r=0.90$, 95% CI $[0.86,0.93]$, $p<0.001$), indicating high agreement between our manual and the Azerion Ad Expert ad counts. The ad counts differed by more than one ad in only three website–ad-blocker observations, likely due to screenshots not always capturing the full webpage. These results show that Azerion Ad Expert accurately counts ads on webpages.
@@
cross-referenced that third-party domain with WhoTracks.me’s database [^8] https://whotracks.me of known trackers (e.g., Adobe Audience Manager tracker uses a demdex.net domain; [49]). By subtracting the baseline number of trackers on a website from the measured number with the ad blocker, we obtained the measure of the number of blocked trackers.
@@ 6.1 Descriptive Results for Ad Blocker Effectiveness
Figure 3: Distribution of the number of blocked ads (Panel A) and trackers (Panel B) per website and browser instance. The red dashed horizontal line at 0 highlights the threshold below which ad-blocking or anti-tracking effectiveness is negative. Multiplying the 9 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3, and MV3+) with the number of websites (114) yields the number of observations (N = 1,026).
Blocked Ads (Panel A): The distributions for the number of blocked ads are similar across ad blockers, indicating comparable ad-blocking effectiveness. MV2 ad blockers (Adblock Plus MV2, AdGuard MV2, Stands MV2, and uBlock MV2) show medians ranging from about 2.15 to 2.55 blocked ads. Their MV3 counterparts exhibit similar medians—with the exception of Adblock Plus MV3, showing a lower median (around 2.10) and a minimum of –0.7. We obtain negative values when a website displays more ads with the ad blocker than without it (e.g., 12 ads with the ad blocker versus 10 without the ad blocker yields –2 ads per website and browser instance). All ad blocker groups, including the MV3+ ad blocker group, consistently show an upper outlier at 13 blocked ads, indicating that certain websites repeatedly display a high number of ads.
Blocked Trackers (Panel B): The anti-tracking effectiveness varies among ad blockers. Adblock Plus’s MV3 and MV2 instances exhibit medians of about 17—17.8 blocked trackers, while AdGuard’s MV3 and MV2 instances yield medians of about 24 blocked trackers. In contrast, the uBlock’s MV3 and MV2 instances and the MV3+ ad blocker group show higher medians, around 34 blocked trackers. Although most ad blockers display similar central tendencies, some observations fall below 0 blocked trackers—for example, Stands MV2 reaches a minimum of –9 blocked trackers, meaning that the number of blocked trackers sometimes exceeds the baseline. Notably, negative values appear for Stands MV2 but not for Stands MV3. Moreover, the MV3+ ad blocker group achieves a higher median anti-tracking effectiveness than several MV2 ad blockers (Adblock Plus MV2, AdGuard MV2, and Stands MV2) and is similarly higher than most MV3 ad blockers—except for uBlock MV3, whose anti-tracking effectiveness appears to drive the high anti-tracking effectiveness of the MV3+ ad blocker group.
@@
Figure 4: Number of blocked ads and trackers between manifest version 3 (MV3) and manifest version 2 (MV2) ad blocker groups. This figure shows independent t-test comparisons between the mean values of two ad blocker groups, namely MV3 and MV2, regarding the number of blocked ads (Panel A) and trackers (Panel B). The MV3 ad blocker group includes Adblock Plus MV3, AdGuard MV3, Stands MV3, and uBlock MV3 ad blockers, while the MV2 ad blocker group includes Adblock Plus MV2, AdGuard MV2, Stands MV2, and uBlock MV2 ad blockers. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (114) yields the number of observations (N = 912). Error bars represent ±$\pm 1$ SEs.
Blocked Ads (Panel A): Panel A of Figure 4 shows that there is no significant difference in ad-blocking effectiveness between the MV3 and MV2 ad blocker groups ($\Delta=-0.04$, $p=0.92$). The mean number of blocked ads is about the same for both groups, and the error bars indicate similar variability within each group.
Blocked Trackers (Panel B): Similarly, Panel B of Figure 4 illustrates no significant difference in anti-tracking effectiveness between the MV3 and MV2 ad blocker groups ($\Delta=1.62$, $p=0.54$). The slight difference in the mean number of blocked trackers is not statistically significant, and the variability, as indicated by the error bars, is comparable between the groups.
@@
Figure 5: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (114) yields the number of observations (N = 912). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (114) yields the number of observations (N = 570).
@@
Likewise, in Panel B, comparisons between the MV3+ ad blocker group and the MV3 ad blockers yield no significant differences in their ad-blocking effectiveness.
Blocked Trackers (Panels C and D): Panel C, however, shows higher anti-tracking effectiveness for Stands MV3 compared to its MV2 instance ($\Delta=6.11$, $p=0.03$). The mean number of blocked trackers for Stands MV2 is about 20.0, while for Stands MV3 it is about 26.1. This increase of 6.11 blocked trackers represents a 30.5% improvement in anti-tracking effectiveness for Stands MV3: $(\frac{26.1-20.0}{20.0})\times 100\%=30.5\%$.
In Panel D, the comparison between the MV3+ ad blocker group and the MV3 ad blockers reveals higher anti-tracking effectiveness of the MV3+ ad blocker group compared to Adblock Plus MV3 ($\Delta=12.66$, $p=0.00$), AdGuard MV3 ($\Delta=6.73$, $p=0.02$), and Stands MV3 ($\Delta=7.03$, $p=0.02$). These results demonstrate that using multiple MV3 ad blockers simultaneously (MV3+) enhances anti-tracking effectiveness by about 61.8% over Adblock Plus MV3 [$(\frac{12.7}{20.5})\times 100\%=61.8\%$], 25.4% over AdGuard MV3 [$(\frac{6.73}{26.4})\times 100\%=25.4\%$], and 26.9% over Stands MV3 [$(\frac{7.03}{26.1})\times 100\%=26.9\%$].
@@ 6.4 Insights from Robustness Tests
Alternative Measures for Ad Blocker Effectiveness The original measures of ad blocker effectiveness might fail to capture certain types of ads or lesser known trackers; need for an alternative measures. Results remain consistent using HTML file size (kb) of loaded websites and number of blocked third-party domains, confirming the robustness of our main results. B.2
@@
Visual Inspection of MV3 and MV2 Ad Blocker Screenshots Potential ad flickering, unintended blockage of legitimate content or cosmetic differences might not be fully captured by our metrics. We detected no noticeable ad flickering or website breakage in screenshots; we observed some differences in website cosmetic placeholders.
@@
The first robustness test (see Appendix B.1) examined whether our findings hold despite variations in sample size and website selection. We repeated our analyses on two separate stratified samples. These samples were stratified by website employee count and Tranco popularity rank, creating samples more representative of the overall population of websites in resources and popularity. The results remained consistent to those of our main sample, confirming the robustness of our findings.
@@
The second robustness test (see Appendix B.2) uses alternative measures of ad blocker effectiveness, namely HTML file size of loaded websites and the number of blocked third-party domains. We measured the HTML file size (kb) of loaded websites as an alternative measure for ad-blocking and anti-tracking effectiveness of ad blockers. T-tests between the MV3 and MV2 instances of ad blockers increase confidence in our main results by showing mostly non-significant differences between manifest versions. An exception is the Stands MV3 ad blocker, which causes significantly lower HTML file sizes than its MV2 counterpart, in line with previous findings. Similarly, the number of blocked third-party domains is an alternative measure for anti-tracking effectiveness to the number of blocked trackers. While the number of blocked trackers is a measure derived from third-party domains cross-referenced with the WhoTracks.me database, sensitive to well-known trackers, the number of blocked third-party domains counts all distinct third-party domains blocked, offering a broader view of the ad blockers’ anti-tracking effectiveness. Despite possible variations due to this more granular measure, MV3 and MV2 ad blocker groups displayed the same anti-tracking effectiveness as before. Thus, the results of this robustness test remain consistent with our findings.
@@
The third robustness test (see Appendix B.3) compared the first versions of MV3 ad blockers (AdGuard MV3 version 0.3.9 and uBlock MV3 version 0.1.22.806) with their MV2 counterparts. This comparison aimed to assess changes in their effectiveness over time, exploring whether ad blocker providers adapted to MV3’s restrictions. Interestingly, the early MV3 versions were equally effective as their later versions, suggesting that the effectiveness of MV3 ad blockers has remained mostly consistent over time. This finding suggests that when ad blocker providers first introduced their MV3 instances of ad blockers, they were likely confident in their effectiveness. The core functionality of these ad blockers did not fluctuate significantly across subsequent releases.
@@
The fourth robustness test (see Appendix B.4) evaluated ad blocker effectiveness on the Firefox browser to address concerns that our Chrome results might be browser-specific. We conducted this robustness test with uBlock as it is currently the only ad blocker on Firefox with a MV3 and MV2 instance. As in our browser-based experiment, we developed a custom HTTP Recorder extension to count the number of blocked trackers, but counted the number of blocked ads per website manually—as Azerion Ad Expert is not available for Firefox browser. We found that our results on the Firefox browser align with those on Chrome browser, indicating that uBlock’s effectiveness remained consistent across browsers and manifest versions. This result indicates that our findings are applicable across the two browsers.
@@
Lastly, we conducted a fifth robustness test by visually inspecting the captured screenshots during the crawling process. These screenshots were taken as the browser scrolled through each webpage, allowing us to compare a subset of MV3 and corresponding MV2 browser instances (100 each) side by side. Our goal was to detect additional ad flickering, unintended blockage of legitimate content (breakage), or cosmetic differences. We expected any ad flickering to appear as discrepancies in ad counts visible in the screenshots but not reflected in our measured ad counts. However, we did not detect noteworthy ad flickering or breakage differences across manifest versions. We observed one instance where an ad was visible under the MV3 ad blocker but absent from our counts, along with five breakage incidents: four where the MV2 ad blocker mistakenly blocked legitimate content allowed by the MV3 ad blocker, and one where the MV3 ad blocker blocked legitimate content that the MV2 ad blocker did not. Despite the general consistency, our visual inspection revealed some cosmetic differences in the form of visible placeholders. In 14 out of 100 comparisons, placeholders were apparent under MV3 but not under MV2, whereas the opposite was true in 2 out of 100 instances. Thus, we conclude that the MV3 update does not introduce notable ad flickering or website breakage, though MV3 ad blockers occasionally produce a less visually appealing experience than their MV2 counterparts.

### v2

@@ 5.3 Selection of Websites for the Browser-Based Experiment
For the main sample of our browser-based experiment, we selected ad-supported websites to observe the impact of ad blockers and evaluate their effectiveness. For this purpose, we utilized a curated list from BuiltWith [^6] https://builtwith.com, which comprised data of websites relying on the Prebid framework. From the list of websites provided by BuiltWith on January 9, 2024, we selected the top 1,000 most popular websites according to their Tranco popularity rank [60]. The Tranco popularity rank of these websites varied from 17 to 17,590. We selected these websites for their reliance on ad-supported content and their use of the Prebid framework, guaranteeing compatibility with the Azerion Ad Expert extension (see Section 5.2). This compatibility is crucial for ensuring the extension accurately records the number of ads, especially for the baseline browser instance using no ad blocker.
Across five separate measurement runs, we consistently obtained results for 924 of these websites, excluding those that redirected users, deployed ad-blocking prevention or anti-automation mechanisms, or were inaccessible. The final list for our main sample represents 924 popular, ad-supported websites across 23 countries.
While our main sample targets popular, ad-supported websites, ad blockers may behave differently on less popular websites or websites with limited resources—for example, because extensions sometimes need to drop less frequently used blocking rules when they hit internal limits on total rule counts. To address this point and improve the external validity of our results, we conducted the browser-based experiment on two additional stratified samples for robustness tests. Using the same BuiltWith dataset, we stratified:
  • One sample by website employee count, yielding 191 websites (including many with zero recorded employees), and
  • Another by Tranco popularity rank, yielding 185 websites (extending as far down as ranks in the millions).
These stratified samples therefore cover both ends of the popularity and resource-availability spectrum, ensuring our findings are not driven solely by high-traffic, well-resourced websites. The resulting diversity in organizational size and popularity enhances the generalizability of our conclusions (see Section 6.4 and Appendix B.1).
@@ 5.4 Definition of Two Metrics for Ad Blocker Effectiveness
For the first metric, the number of blocked ads, we used the extension Azerion Ad Expert. As mentioned (see Section 5.2), this tool recorded the number of display ad slots on a website. By subtracting the measured number of ads on the website while using the ad blocker from the baseline number of ads on the website without using the ad blocker, we obtained the number of blocked ads.
To validate the accuracy of these ad counts, we manually counted ads in collected screenshots from a random subsample of 100 website-ad blocker observations in one of our early runs. A Pearson correlation analysis showed a strong positive correlation ($r=0.90$, 95% CI $[0.86,0.93]$, $p<0.001$), indicating high agreement between our manual and the Azerion Ad Expert ad counts. The ad counts differed by more than one ad in only three website-ad blocker observations, likely due to screenshots not always capturing the full webpage. These results show that Azerion Ad Expert accurately counts ads on webpages.
@@
cross-referenced that third-party domain with WhoTracks.me’s database [^7] https://whotracks.me of known trackers (e.g., Adobe Audience Manager tracker uses a demdex.net domain; [61]). By subtracting the measured number of trackers on the website while using the ad blocker from the baseline number of trackers on the website without using the ad blocker, we obtained the number of blocked trackers.
@@ 6.1 Descriptive Results for Ad Blocker Effectiveness
Figure 3: Distribution of the number of blocked ads (Panel A) and trackers (Panel B) per website and browser instance. The red dashed horizontal line at 0 highlights the threshold below which ad-blocking or anti-tracking effectiveness is negative. Multiplying the 9 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3, and MV3+) with the number of websites (924) yields the number of observations (N = 8,316). Values shown are per-website averages across five runs.
Blocked Ads (Panel A): The distributions for the number of blocked ads are similar across ad blockers, indicating comparable ad-blocking effectiveness. MV2 ad blockers (Adblock Plus MV2, AdGuard MV2, Stands MV2, and uBlock MV2) show medians ranging from about 0.4 to 0.6 blocked ads. Their MV3 counterparts exhibit similar medians—with the exception of Adblock Plus MV3, showing a lower median (around 0.2) and a minimum of –5. We obtain negative values when a website displays more ads with the ad blocker than without it (e.g., 12 ads with the ad blocker versus 10 without the ad blocker yields –2 ads per website and browser instance). All ad blocker groups, including the MV3+ ad blocker group, consistently show an upper outlier at 13 blocked ads, indicating that certain websites repeatedly display a high number of ads.
Blocked Trackers (Panel B): The anti-tracking effectiveness varies among ad blockers. Adblock Plus’s MV3 and MV2 instances exhibit medians of about 8–10.6 blocked trackers, while AdGuard’s MV3 and MV2 instances yield medians of about 17 blocked trackers. In contrast, the uBlock’s MV3 and MV2 instances and the MV3+ ad blocker group show higher medians, around 22 blocked trackers. Although most ad blockers display similar central tendencies, some observations fall below 0 blocked trackers—for example, Adblock Plus MV2 reaches a minimum of –27 blocked trackers, meaning that the number of blocked trackers sometimes exceeds the baseline. Notably, negative values also appear for other instances (e.g., down to –20.6 for Adblock Plus MV3 and –1.8 for Stands MV3). Moreover, the MV3+ ad blocker group achieves a higher median anti-tracking effectiveness than several MV2 ad blockers (Adblock Plus MV2, AdGuard MV2, and Stands MV2) and is similarly higher than most MV3 ad blockers. The anti-tracking effectiveness of uBlock MV3 appears to drive these results in the MV3+ ad blocker group, suggesting that multiple MV3 ad blockers do not hinder each other through the global rule limit.
@@
Figure 4: Number of blocked ads and trackers between manifest version 3 (MV3) and manifest version 2 (MV2) ad blocker groups. This figure shows independent t-test comparisons between the mean values of two ad blocker groups, namely MV3 and MV2, regarding the number of blocked ads (Panel A) and trackers (Panel B). The MV3 ad blocker group includes Adblock Plus MV3, AdGuard MV3, Stands MV3, and uBlock MV3 ad blockers, while the MV2 ad blocker group includes Adblock Plus MV2, AdGuard MV2, Stands MV2, and uBlock MV2 ad blockers. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (924) yields the number of observations (N = 7,392). Values shown are per-website averages across five runs; error bars represent $\pm 1$ SE across websites.
Blocked Ads (Panel A): Panel A of Figure 4 shows that there is no significant difference in ad-blocking effectiveness between the MV3 and MV2 ad blocker groups ($\Delta=-0.04$, $p=0.41$). The mean number of blocked ads is about the same for both groups, and the error bars indicate similar variability within each group.
Blocked Trackers (Panel B): Panel B of Figure 4 illustrates a significant difference in anti-tracking effectiveness between the MV3 and MV2 ad blocker groups ($\Delta=1.78$, $p<0.001$). The difference of about 2 blocked trackers in the mean number of blocked trackers is statistically significant, and the variability, as indicated by the error bars, is comparable between the groups.
@@
Figure 5: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (924) yields the number of observations (N = 7,392). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (924) yields the number of observations (N = 4,620). Values shown are per-website averages across five runs; error bars represent $\pm 1$ SE across websites.
@@
In Panel B, comparisons between the MV3+ ad blocker group and the MV3 ad blockers show a significant difference for Adblock Plus ($\Delta=0.28$, $p<0.01$), while AdGuard, Stands, and uBlock show no significant differences.
Blocked Trackers (Panels C and D): Panel C shows higher anti-tracking effectiveness for Stands MV3 compared to its MV2 instance ($\Delta=5.19$, $p<0.001$). The mean number of blocked trackers for Stands MV2 is about 11.3, while for Stands MV3 it is about 16.5. This increase of 5.19 blocked trackers represents a 45.9% improvement for Stands MV3: $(\frac{16.5-11.3}{11.3})\times 100\%=45.9\%$. In addition, Adblock Plus MV3 also exceeds its MV2 instance ($\Delta=1.91$, $p<0.001$), representing a 21.5% improvement for Adblock Plus MV3: $(\frac{10.8-8.89}{8.89})\times 100\%=21.5\%$. AdGuard and uBlock show no significant differences.
In Panel D, the comparison between the MV3+ ad blocker group and the MV3 ad blockers reveals higher anti-tracking effectiveness of the MV3+ ad blocker group compared to Adblock Plus MV3 ($\Delta=10.29$, $p<0.001$), AdGuard MV3 ($\Delta=4.44$, $p<0.001$), and Stands MV3 ($\Delta=4.58$, $p<0.001$). These results demonstrate that using multiple MV3 ad blockers simultaneously (MV3+) enhances anti-tracking effectiveness by about 95.3% over Adblock Plus MV3 $(\frac{10.29}{10.8})\times 100\%=95.3\%$], 26.6% over AdGuard MV3 $(\frac{4.44}{16.7})\times 100\%=26.6\%$], and 27.8% over Stands MV3 $(\frac{4.58}{16.5})\times 100\%=27.8\%$].
@@ 6.4 Insights from Robustness Tests
Alternative Measures for Ad Blocker Effectiveness The original measures of ad blocker effectiveness might fail to capture certain types of ads or lesser-known trackers; need for alternative measures. Results remain consistent using HTML file size (kb) of loaded websites and number of blocked third-party domains, confirming the robustness of our main results. B.2
@@
Visual Inspection of MV3 and MV2 Ad Blocker Screenshots Potential ad flickering, unintended blockage of legitimate content or cosmetic differences might not be fully captured by our metrics. We detected no noticeable ad flickering or website breakage in screenshots; we observed some differences in the visibility of cosmetic placeholders. B.5
@@
The first robustness test (see Appendix B.1) examined whether our findings hold despite variations in sample size and website selection. We repeated our analysis on two separate stratified samples. These samples were stratified by website employee count, since websites with varying resources might serve ads differently, and Tranco popularity rank, creating samples more representative of the overall population of websites in resources and popularity. In this test, the Tranco popularity rank of tested websites varied from 158 to more than one million, and the number of employees from 0 to 10,000, covering a wide range of website types. Specifically, in the sample stratified by website employee count (191 websites), MV3 and MV2 did not differ in blocked ads; for trackers, only Stands improved under MV3 ($\Delta=7.03$, $p<0.001$). In the sample stratified by Tranco popularity rank (185 websites), MV3 and MV2 did not differ in either metric. In both stratified samples, the MV3+ group blocked significantly more trackers than Adblock Plus MV3, AdGuard MV3, and Stands MV3 (no significant difference vs. uBlock MV3). Taken together, these stratified-sample results are largely consistent with those of our main sample, reinforcing the robustness of our main findings.
@@
The second robustness test (see Appendix B.2) used alternative measures of ad blocker effectiveness, namely HTML file size of loaded websites and the number of blocked third-party domains. We measured the HTML file size (kb) of loaded websites as an alternative measure for ad-blocking and anti-tracking effectiveness of ad blockers. T-tests between the MV3 and MV2 instances of ad blockers increase confidence in our main results by showing mostly non-significant differences between manifest versions. An exception is the Stands MV3 ad blocker, which caused significantly lower HTML file sizes than its MV2 counterpart, in line with previous findings. Similarly, the number of blocked third-party domains is an alternative measure for anti-tracking effectiveness to the number of blocked trackers. While the number of blocked trackers is a measure derived from third-party domains cross-referenced with the WhoTracks.me database, sensitive to well-known trackers, the number of blocked third-party domains counts all distinct third-party domains blocked, offering a broader view of the ad blockers’ anti-tracking effectiveness. Despite possible variations due to this more granular measure, the pattern mirrored our tracker results at the individual level: Adblock Plus and Stands blocked more domains under MV3, while AdGuard and uBlock showed parity between MV3 and MV2. Overall, these alternative measures reinforce our main findings.
@@
The third robustness test (see Appendix B.3) compared the first versions of MV3 ad blockers (AdGuard MV3 version 0.3.9 and uBlock MV3 version 0.1.22.806) with their MV2 counterparts. This comparison aimed to assess changes in their effectiveness over time, exploring whether ad blocker providers adapted to MV3’s restrictions. Interestingly, the early MV3 versions were equally effective as their later versions, suggesting that the effectiveness of MV3 ad blockers has remained mostly consistent over time. This finding suggests that when ad blocker providers first introduced their MV3 instances of ad blockers, they were likely confident in their effectiveness. The core functionality of these two ad blockers did not fluctuate significantly across subsequent releases.
@@
The fourth robustness test (see Appendix B.4) evaluated ad blocker effectiveness on the Firefox browser to address concerns that our Chrome results might be browser-specific. We conducted this robustness test with uBlock as it is currently the only ad blocker on Firefox with a MV3 and MV2 instance. As in our browser-based experiment, we developed a custom HTTP Recorder extension to count the number of blocked trackers, and counted the HTML file size of loaded websites—as Azerion Ad Expert is not available for Firefox browser. We found that our results on the Firefox browser align with those on Chrome browser, indicating that uBlock’s effectiveness remained consistent across browsers and manifest versions. This result indicates that our findings are applicable across the two browsers.
@@
Lastly, we conducted a fifth robustness test by visually inspecting the captured screenshots during the crawling process. These screenshots were taken as the browser scrolled through each webpage, allowing us to compare a subset of the tested MV3 and corresponding MV2 browser instances (100 each) side by side. Our goal was to detect additional ad flickering (the visibility of ads, that were not present in our ad counts), unintended blockage of legitimate content (content that is considered regular website functionality irrespective of ads), or cosmetic placeholders (the visibility of empty placeholders on webpages). Two evaluators independently evaluated outcomes based on a provided set of questions, one per outcome. We report results as Evaluator 1 (Evaluator 2). The two evaluators found no substantial ad-flickering discrepancies between MV3 and MV2: 4 (4) missed ads, split evenly across MV3 and MV2. They also observed no major breakage differences: 5 (5) MV3-only versus 5 (3) MV2-only cases. However, they found that MV3 left over cosmetic placeholders in 21% (22%) of cases versus 0% (0%) under MV2. Further details on how we conducted this test and examples can be found in Appendix B.5. In summary, we conclude that the MV3 update did not introduce notable additional ad flickering or website breakage, though MV3 ad blockers occasionally produced a less visually appealing experience than their MV2 counterparts due to the increased visibility of cosmetic placeholders.